Test of DX LGSW with CalUnit and rigid secondary (afternoon)

Problems with Arbitrator
  • Restart not possible in situation of shutdown of LAS and powering_up of LGSW. During transition state no command accepted on LGSW side and timed out on Arbitrator side.
  • Previous situation happened due to blocking of Arbitrator (graying out). This happened several times.
After setting all up red laser could be found left of yellow.With following parameters the beams could be caught (very much trial and error and not representative)

Right PSF 2ndary
X Y Z Rx Ry Rz
-3 -6 -5.3 60 -200 0
Right PSF tertiary
Tip Tilt Piston Rz
100 -1000 0 0

Setup Lasers for heater test (evening)

With setting up the lasers for the heater test Laser3 on DX was not working properly. After closing and opening internal shutter with help of ranger (SOFTWARE_KEY_POSITION) issue was resolved.

LgswUnitSx calibration with AdSec

PBL with AdSec start. The ADSEC.L.OFLMODES variable is set as a string at startup while it should be real. A workaround is to delete it and to restart the fast diagnostic service.

We start acquiring new IMs having fixed the slope computing bug.

Pupil motors are:

Beam X Y
Blue 325 421
Yellow 376 438
Red 487 432

New jitter initial position: 20151211_155637

# modes IM REC IM LGSW REC
10 - 20151211_154845 20151211_155000 -
10 20151211_155000 20151211_155258 20151211_155400
50 20151211_155400 20151211_155956 20151211_160100
150 20151211_160100 20151211_161154 20151211_161400

We find out that we are acquiring IMs at 1kHz, we reduce loop speed at 610Hz:

Beam X Y
Blue 325 411
Yellow 376 438
Red 487 432

# modes IM REC IM LGSW REC
150 20151211_161400 20151211_161933 20151211_162200

We define new modal amplitudes having normalized all slopes rms to 0.05: modal amplitude: 20151211_162400_rms0.05

On this modal amplitude we define several modal histories:

# modes Modal history
150 20151211_163400_150modes
50 20151211_163500_50modes
10 20151211_163600_10modes

We recenter again pupils and acquire a new jitter initial position: 20151211_164257. There is no way to have all the 3 pupils unvignetted frown, sad smile

# modes IM REC IM LGSW REC
150 20151211_162200 20151211_164434 20151211_164600

We see a sort of ear at the lower right edge of the blue pupil. It is much more visible in mode 1 Y slopes.

  • Modal amplitude: 20151211_164800_rms0.10
  • Modal amplitude: 20151211_164900_rms0.20

From these 2 modal amplitudes we create histories:

# modes Modal amplitude Modal history
150 20151211_164800_rms0.10.fits 20151211_165000_150modes
50 20151211_165100_50modes
10 20151211_165200_10modes

Using MH 20151211_165000_150modes:

# modes IM REC IM LGSW REC
150 20151211_164600 20151211_165247 20151211_165300
150 20151211_164600 20151211_165912 ears are visible in the TT modes

We try to increase the TT gain during IM acquisition from 0.05 to 0.2:

# modes IM REC IM LGSW REC
150 20151211_164600 20151211_170128 ears...

We fear that the pupil can be misaligned. Up to now we aligned them by putting 0 flux on the central subaperture.

Beam X Y
Blue 325 431
Yellow 366 448
Red 497 432

New pupil target tag: 20151211_171015 and jitter initial position: 20151211_171324

# modes IM REC IM LGSW REC COMBO REC
150 20151211_164600 20151211_171626 20151211_171800 20151211_172100

The last LGSW only combined rec it is not working fine. The lower right edge of the pupils is always quite instable.

# modes IM REC IM LGSW REC COMBO REC
50 20151211_164600 20151211_171626 20151211_173400 20151211_173500

TN Notes
20151211_173602 50 modes CL, w/o disturb
20151211_173621

# modes IM REC IM LGSW REC COMBO REC
150 20151211_173400 20151211_173939 20151211_174100 20151211_174200

This one seems to work fine...

TN Notes
20151211_174325 150 modes CL, w/o disturb
20151211_174359
20151211_174602 1" disturb, gains= 0.5, 600Hz
20151211_174624
20151211_174643
20151211_175028 1kHz
20151211_175059
20151211_175652 0.8" disturb, gs= [0.5, 0.4, 0.4], 1kHz
20151211_175708
20151211_175725

LgswUnitSx calibrations for TCoG48 algorithm

We define a new subaperture set based on TCoG48 slope computing algorithm: 20151211_181800

# modes IM REC IM LGSW REC COMBO REC
10 20151211_174100 20151211_182550 20151211_182700 -
50 20151211_182700 20151211_182841 20151211_183000 -
150 20151211_183000 20151211_183209 20151211_183300 20151211_183400

TN Notes
20151211_183720 150 modes CL, 0.8" disturb
20151211_183745
20151211_183759

HEXA 4.67 -0.137 7.911 234 -93

Beam X Y
Blue 325 431
Yellow 366 448
Red 497 432

Pupil target tag: 20151211_ 184842

Jitter initial position: 20151211_171324

We close the loop on 150 modes, LGSW only. The best combination of gains seems to be [0.5, 0.4, 0.4]. We load 1.0" disturb:

TN Notes
20151211_190535 150 modes CL, 1.0" disturb
20151211_190557
20151211_192827
20151211_193014

Now we start the FLAO, we feed it with the pennina. We center the pyramid board to -71.95, -46.7, 35.0 mm. Pupil rerotator is at 88.4deg, instrument rotator at 340.98deg. Pyr configured as BIN2, 975Hz.

We build combined recs for FLAO:

REC TAG LGSW REC FLAO REC COMBO REC
TCoG48, BIN2 20151211_183400 20151202_213530 20151211_202800
TCoG48, BIN4 20151204_191745 20151211_202900

14:00: we check the best focus for the N3.75 camera. Z stage is at 50mm.

Now we close on LGSW only. Z stage is at 35mm for the best focus position on LGSW...

TN Notes
20151211_222841 doing best focus on N30
20151211_222905

-- WolfgangGaessler - 12 Dec 2015
Topic revision: r3 - 08 Jan 2016, MarcoBonaglia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback